# The FAST-SETI Milky Way Globular Cluster Survey I: A Pilot Multibeam On-the-Fly Search of Five Globular Clusters at L-Band

### v1

@@ 5.1 Technosignature prevalence in metal-rich globular clusters
5.2 MBPS versus prior ON–OFF/MBCM strategies
Classical ON–OFF and multibeam coincidence matching (MBCM) are snapshot tests: a candidate is accepted if it vanishes in an OFF pointing (ON–OFF) or is present in a designated on-beam while absent in others at one epoch (MBCM). Both schemes control RFI but rely on stationarity and a single on-beam geometry, which reduces on-source duty cycle and creates failure modes for intermittency, beam miscentering, or extended fields (tao2022sensitive; luan2023multibeam).
MBPS converts vetting from snapshot heuristics to a deterministic, geometry-driven classification test (Huang2023MBPS). During controlled slews, the multibeam footprint imposes a known time–beam illumination sequence; a sky-localized signal must appear only within its predicted illumination windows, be absent from non-illuminated beams at the same times, and step through beams in the order and cadence set by the scan. This yields orthogonal, in situ checks (spatial exclusivity, simultaneous references, and temporal ordering) without sacrificing duty cycle.
The modification introduced here extends MBPS from point targets to an entire field by (i) defining generalized on–target windows (gOTWs) from the intersection of each beam track with a buffered cluster mask and (ii) enforcing the same timing/ordering logic within each declination stripe (BG–A to BG–D) and across two complementary passes (Sec. 3.2). Passes that do not illuminate a position are not penalized, while doubly illuminated positions must show consistent drift and timing; where applicable, non-coherent combination across crossings improves sensitivity by $\propto\sqrt{N_{\rm cross}}$ (Sec. 3.3). In practice, this framework preserves the array-wide simultaneity veto of MBCM, restores ON–OFF–level rigor without inter-pointing assumptions, and adds a predictive temporal model that typical site or comb RFI cannot mimic. Applied to our data, no event satisfies the full set of MBPS criteria over the illuminated geometry (Sec. 4.2), yielding a robust null under extended-target conditions.
5.3 What we were sensitive to—and what we were not
Our search is optimized for Hz-wide, slowly drifting carriers across 1.05–1.45 GHz, detected within $\sim 60$ s OTWs and, when geometry permits, stacked across available OTB crossings (upper bound $\sqrt{N_{\rm cross}}$ improvement; §3.3). We are intrinsically less sensitive to (i) broadband or rapidly hopped emissions without matched filters, (ii) very low duty-cycle beacons whose emission seldom overlaps an OTW, and (iii) signals outside L-Band. The thresholds in Table 4 therefore function as a link-budget guide: at kiloparsec distances, isotropic transmission is energetically extravagant, whereas beamed systems (large dishes/arrays) can meet the thresholds with far lower transmitter power.
@@ 5.3 What we were sensitive to—and what we were not
Because ${\rm EIRP}=P_{\rm tx}G_{\rm tx}$, our limits can be read as required $P_{\rm tx}$ at assumed $G_{\rm tx}$. For reference, at 1.3 GHz a 300 m dish ($\eta\sim 0.6$) has $G_{\rm tx}\sim 10^{7}$ (70 dBi), while 100 m dishes yield $G_{\rm tx}\sim 10^{6}$ (60 dBi). Thus ${\rm EIRP}_{\min}\sim 10^{16}$ W maps to $P_{\rm tx}\sim 10^{8}\text{--}10^{10}$ W depending on aperture; by comparison, the historic Arecibo S-band radar achieved $\sim 10^{13}$ W EIRP (e.g., Siemion2013Kepler). This framing clarifies where beamed vs. isotropic strategies land relative to our limits.
5.3.1 Linear-polarization statistics and implications for Stokes-parameter filtering
@@
5.3.2 Detection probability for intermittent emitters
During the MBOTF raster, a sky position is illuminated in gOTWs of duration $t_{\rm OTW}\!\approx\!60$ s, separated by $\sim\!115$ s along the five-beam line (Fig. 1). Within each illuminated crossing we analyze spectra with cadence $\Delta t\!\approx\!10.2$ s . Therefore, the single-crossing sensitivity quoted in §3.3 (e.g., $S_{\min}\!\approx\!0.50$ Jy for $t_{\rm OTW}\!\approx\!60$ s) applies to continuous emission within the window.
@@ 5.3.2 Detection probability for intermittent emitters
5.4 Figure of Merit and placement on the CW rate–sensitivity plane
Following the continuous–wave (CW) comparison framework popularized by Enriquez2017BL692, we assess our survey in the plane of sensitivity versus coverage: the abscissa is $\log_{10}(EIRP_{\min}/{\rm W})$ and the ordinate is the transmitter rate, defined as the per–star, per–fractional–bandwidth upper limit $(N_{\star}\,\nu_{\rm rel})^{-1}$. In this space, moving left indicates sensitivity to weaker transmitters (smaller $EIRP_{\min}$) while moving down corresponds to tighter occurrence limits (larger $N_{\star}\nu_{\rm rel}$).
@@ 5.4 Figure of Merit and placement on the CW rate–sensitivity plane
A convenient scalar summary is the CW transmitter figure of merit (CWTFM; Enriquez2017BL692),
@@
Figure 6: Comparison of technosignature survey limits in the plane of transmitter rate versus sensitivity, following the framework of Enriquez2017BL692. The abscissa represents the minimum detectable Equivalent Isotropic Radiated Power ($\mathrm{EIRP}_{\min}$), and the ordinate denotes the upper limit on transmitter rate, defined as $(N_{\star}\nu_{\mathrm{rel}})^{-1}$. The dark cyan diamond marks the constraints established by this work (FAST GC Survey Phase I). Historical and contemporaneous limits are plotted for comparison, with data drawn from: Verschuur1973, Tarter1980, Valdes1986, HorowitzSagan1993, Project Phoenix (Backus2002), Siemion2013Kepler, Harp2016, GrayMooley2017, Price2020BL1327, Gajjar2021GC, tao2022sensitive, luan2023multibeam, Choza2024, Painter2025 and Luan_2025. The gray band traces the characteristic sensitivity–coverage trend of these prior searches. Our survey’s position in the lower-left quadrant highlights its joint optimization of depth (via FAST’s high gain) and breadth (via the high stellar density of globular clusters and wide L-band coverage).
5.5 Conclusions

### v3

@@ V.1 Technosignature prevalence in metal-rich globular clusters
V.2 MBPS versus prior ON–OFF/MBCM strategies
Classical ON–OFF and multibeam coincidence matching (MBCM) are snapshot tests: a candidate is accepted if it vanishes in an OFF pointing (ON–OFF) or is present in a designated on-beam while absent in others at one epoch (MBCM). Both schemes control RFI but rely on stationarity and a single on-beam geometry, which reduces on-source duty cycle and creates failure modes for intermittency, beam miscentering, or extended fields (Tao et al., 2022; Luan et al., 2023).
MBPS converts vetting from snapshot heuristics to a deterministic, geometry-driven classification test (Huang et al., 2023). During controlled slews, the multibeam footprint imposes a known time–beam illumination sequence; a sky-localized signal must appear only within its predicted illumination windows, be absent from non-illuminated beams at the same times, and step through beams in the order and cadence set by the scan. This yields orthogonal, in situ checks (spatial exclusivity, simultaneous references, and temporal ordering) without sacrificing duty cycle.
The modification introduced here extends MBPS from point targets to an entire field by (i) defining generalized on–target windows (gOTWs) from the intersection of each beam track with a buffered cluster mask and (ii) enforcing the same timing/ordering logic within each declination stripe (BG–A to BG–D) and across two complementary passes (Sec. III.2). Passes that do not illuminate a position are not penalized, while doubly illuminated positions must show consistent drift and timing; where applicable, non-coherent combination across crossings improves sensitivity by $\propto\sqrt{N_{\rm cross}}$ (Sec. III.3). In practice, this framework preserves the array-wide simultaneity veto of MBCM, restores ON–OFF–level rigor without inter-pointing assumptions, and adds a predictive temporal model that typical site or comb RFI cannot mimic. Applied to our data, no event satisfies the full set of MBPS criteria over the illuminated geometry (Sec. IV.2), yielding a robust null under extended-target conditions.
V.3 What we were sensitive to—and what we were not
Our search is optimized for Hz-wide, slowly drifting carriers across 1.05–1.45 GHz, detected within $\sim 60$ s OTWs and, when geometry permits, stacked across available OTB crossings (upper bound $\sqrt{N_{\rm cross}}$ improvement; §III.3). We are intrinsically less sensitive to (i) broadband or rapidly hopped emissions without matched filters, (ii) very low duty-cycle beacons whose emission seldom overlaps an OTW, and (iii) signals outside L-Band. The thresholds in Table 4 therefore function as a link-budget guide: at kiloparsec distances, isotropic transmission is energetically extravagant, whereas beamed systems (large dishes/arrays) can meet the thresholds with far lower transmitter power.
@@ V.3 What we were sensitive to—and what we were not
Because ${\rm EIRP}=P_{\rm tx}G_{\rm tx}$, our limits can be read as required $P_{\rm tx}$ at assumed $G_{\rm tx}$. For reference, at 1.3 GHz a 300 m dish ($\eta\sim 0.6$) has $G_{\rm tx}\sim 10^{7}$ (70 dBi), while 100 m dishes yield $G_{\rm tx}\sim 10^{6}$ (60 dBi). Thus ${\rm EIRP}_{\min}\sim 10^{16}$ W maps to $P_{\rm tx}\sim 10^{8}\text{--}10^{10}$ W depending on aperture; by comparison, the historic Arecibo S-band radar achieved $\sim 10^{13}$ W EIRP (e.g., Siemion et al., 2013). This framing clarifies where beamed vs. isotropic strategies land relative to our limits.
V.3.1 Linear-polarization statistics and implications for Stokes-parameter filtering
@@
V.3.2 Detection probability for intermittent emitters
During the MBOTF raster, a sky position is illuminated in gOTWs of duration $t_{\rm OTW}\!\approx\!60$ s, separated by $\sim\!115$ s along the five-beam line (Fig. 1). Within each illuminated crossing we analyze spectra with cadence $\Delta t\!\approx\!10.2$ s . Therefore, the single-crossing sensitivity quoted in §III.3 (e.g., $S_{\min}\!\approx\!0.50$ Jy for $t_{\rm OTW}\!\approx\!60$ s) applies to continuous emission within the window.
@@ V.3.2 Detection probability for intermittent emitters
V.4 Figure of Merit and placement on the CW rate–sensitivity plane
Following the continuous–wave (CW) comparison framework popularized by Enriquez et al. (2017), we assess our survey in the plane of sensitivity versus coverage: the abscissa is $\log_{10}(EIRP_{\min}/{\rm W})$ and the ordinate is the transmitter rate, defined as the per–star, per–fractional–bandwidth upper limit $(N_{\star}\,\nu_{\rm rel})^{-1}$. In this space, moving left indicates sensitivity to weaker transmitters (smaller $EIRP_{\min}$) while moving down corresponds to tighter occurrence limits (larger $N_{\star}\nu_{\rm rel}$).
@@ V.4 Figure of Merit and placement on the CW rate–sensitivity plane
A convenient scalar summary is the CW transmitter figure of merit (CWTFM; Enriquez et al., 2017),
@@
Figure 6: Comparison of technosignature survey limits in the plane of transmitter rate versus sensitivity, following the framework of Enriquez et al. (2017). The abscissa represents the minimum detectable Equivalent Isotropic Radiated Power ($\mathrm{EIRP}_{\min}$), and the ordinate denotes the upper limit on transmitter rate, defined as $(N_{\star}\nu_{\mathrm{rel}})^{-1}$. The dark cyan diamond marks the constraints established by this work (FAST GC Survey Phase I). Historical and contemporaneous limits are plotted for comparison, with data drawn from: Verschuur (1973), Tarter et al. (1980), Valdes and Freitas (1986), Horowitz and Sagan (1993), Project Phoenix (Backus et al., 2002), Siemion et al. (2013), Harp et al. (2016), Gray and Mooley (2017), Price et al. (2020), Gajjar et al. (2021), Tao et al. (2022), Luan et al. (2023), Choza et al. (2024), Painter et al. (2025) and Luan et al. (2025). The gray band traces the characteristic sensitivity–coverage trend of these prior searches. Our survey’s position in the lower-left quadrant highlights its joint optimization of depth (via FAST’s high gain) and breadth (via the high stellar density of globular clusters and wide L-band coverage).
V.5 Conclusions